OpenStudy (anonymous):

The trains A and B are traveling in the same direction on parallel tracks. Train A is traveling at 100 mph and train Bis traveling at 120 mph. Train A passes a station at 11:20 pm. If train B passes the same station at 11:35 at what time will train B catch up to train A ?

OpenStudy (anonymous):

let Sb = distance travelled by train B and Sa distance travelled by A - this is when they pass. then Sb = Sa + 25 (25 miles - distance travelled by A in 15 minutes) let t be the time then Sb = 120t and Sa = 100t substituting: 120t = 25 + 100t 20t = 25 t = 1.25 hours so the time when they pass is 11.35 + 1.15 = 12.50

OpenStudy (anonymous):

Thanks !!!! :)

OpenStudy (anonymous):

no probs